DX Field Data - 20140422 UT

Observers: DMiller (Tucson), JStorm, IIlyn (LBTO)
Telescope Operator: GBechetti (LBTO)
AO Support:: DMiller (Tucson)
Telescope Support:: MMidkiff (LBTO)
Instrument Support:: ESolheid (LBTO)
Engineering Support: DAshby (Tucson)



09:00 Switch to PEPSIPFU on BOTH

09:30 Trouble connection to AGW7. Called Taras. Config file AGW7.cfg had been changed for testing with AGW4. Taras changed it back.

10:30 Successful Binocular preset in ACQUIRE mode. Geno find stars on both side in about 5 minutes

10:40 Successful Binocular preset in ACTIVE mode. Collimate in a few iterations. Seeing on GUIDER and DIMM ~ 0.8"

11:00 See the LEFT hotspot jump a couple times. The WFSC images were elongated. Geno added a note to IT 5154

11:20 Changes the RIGHT WFScam_hotspot_x, y from 170.6, 245 to 172.0, 253.6 to center the pupil (was one subap too high). Not clear why the pupil position would change.

11:30 Ilya not able to send preset. IIF complains that target object on set. Will work on it tomorrow.

11:45 J and I collect through focus images on guider and wfsc to check optical axis alignment.
Topic revision: r1 - 22 Apr 2014, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback